# Breaking Language Barriers: Cross-Lingual Continual Pre-Training at Scale

### v1

@@ 1 Introduction
Our main focus lies in the comparative analysis between the two strategies, including their scaling behaviors, the robustness of scaling, and their corresponding impact factors. For this purpose, we fit a new extended scaling law for CPT, derived from Hoffmann et al. (2022). Our findings are outlined as follows:
@@ 2.1 Training Framework
To simplify the experiments, we use identical training strategies for all three pre-training sets. All models are pre-trained with a context length of 2048 and undergo training on tokens equivalent to 20 times the model size (e.g., a 5B model is trained on 100B tokens). Although this is far from the extensive pre-training seen in recent practices  Touvron et al. (2023), as outlined in Hoffmann et al. (2022), the 20x trained token count is sufficient to demonstrate the loss-data scaling relationship. Our learning rate (LR) schedule features a cosine LR decay from a maximum LR of $2\times 10^{-4}$ and an LR warm-up, which increases the LR to the maximum in the first 5% of the training session. We use different batch sizes for different parameter sizes, as shown in Table 1.
@@ 2.2 Model and Data
Our English training data is primarily sampled from the RedPajama dataset Computer (2023), while the Chinese training data was acquired from the public web, undergoing filtering and deduplication processes. To study langauge robustness of the CPT strategy, we also conduct experiments on other languages, including French and Russian. We take their corresponding subsets from mC4 Raffel et al. (2019) as pre-training data. An total of $10^{6}$ tokens are held out from each respective training set as validation sets, remaining consistent across different models.
@@ 4.4 CPT Models Generalize to Downstream Tasks
We assessed the models on their respective language splits of multi-lingual benchmarks to ensure fair comparison. The results indicate that all three languages show improved performance compared to the original pre-trained model, demonstrating that CPT enhances benchmark performance across different languages and scenarios.
@@ 5.1 What is the Compute-Optimal Allocation between Parameter Size and Data?
According to the framework established in Section 3, we can determine the optimal allocation between model parameters $N_{opt}$ and training data $D_{opt}$ by minimizing the predicted loss $L$ with respect to data $D$ and parameter size $N$. More specifically, by optimizing Equation 2, we estimate the optimal training data and model parameters for pre-training from scratch to be:
@@
A visualization of the efficient frontier of model parameter $N$ with respect to compute over the IsoLoss contour is shown in Figure 4. We find that the optimal parameters for continual pre-training differ from those for pre-training from scratch, favoring less compute for the same model sizes. This aligns with the nature of cross-lingual transfer learning, where the model in continual pre-training is "pre-matured" due to prior knowledge acquired in the source language. This suggests that, in continual pre-training, using a larger language model is preferred over pre-training on a larger dataset.
@@ 5.2 Does Replaying from Source Language Prevent Catastrophic Forgetting?
To prevent catastrophic forgetting of the original distributions during continual pre-training, we investigate methods that replay data from the source language during pre-training. We use the term replaying to refer to the practice of mixing data from the source language during continual pre-training on the target language.
@@ 6 Related Work
Scaling laws help us understand how model performance changes with the size of the model and the amount of data. Kaplan et al. (2020) first introduced a detailed scaling law for large language models, demonstrating a clear relationship between model size, training data, and performance. Hoffmann et al. (2022) further explored this by emphasizing the trade-off between model size and data quantity, suggesting a compute-optimal allocation of data and parameters. Recent studies have examined scaling laws under specific conditions. Hernandez et al. (2022) and Muennighoff et al. (2023) focused on the diminishing returns from repeated tokens and excessive parameters. Tay et al. (2022) and Frantar et al. (2023) investigated how different model architectures impact scaling. Scaling laws are also relevant in the context of newer pre-training methods, such as parameter-efficient fine-tuning (PEFT) Kalajdzievski (2024) and Mixture-of-Experts (MoE) Krajewski et al. (2024).
@@
Transfer learning aims to enhance performance on new tasks by adapting pre-trained models with out-of-domain data. This process is more efficient when the source and target domains are closely related Pan and Yang (2009); Zhuang et al. (2020). Cross-lingual pre-training leverages language-independent knowledge embedded in pre-trained LLMs to improve performance in the target language Wu et al. (2019); Yosinski et al. (2014). Transfer learning is often studied within the context of limited-scale post-training, but it has been shown to be effective at a large pre-training scale with proper techniques Gupta et al. (2023). A significant challenge in transfer learning is catastrophic forgetting Winata et al. (2023), where the model’s ability in the original training domain degrades during transfer learning. Various strategies have been proposed to mitigate catastrophic forgetting, including modified learning rate schedules Ibrahim et al. (2024); Gupta et al. (2023); Winata et al. (2023), data replay Ostapenko et al. (2022), and regularization Farajtabar et al. (2020). Our work combines data replay and modified learning rate schedules to combat catastrophic forgetting.
@@ 7 Conclusion
In this study, we utilized publicly accessible datasets for pre-training. Although the Chinese dataset and mC4 dataset attempt to clean and create language-specific training splits, they cannot entirely prevent the contamination of English at a more granular level. This is particularly challenging due to the inherent nature of many languages, such as French, which often incorporate English words. To estimate the computational effort for different languages, we counted the number of samples processed in each language training split. This approach may be imprecise if the dataset contains a large amount of text in other languages. This issue highlights the need for future research to conduct a more in-depth analysis of the impact of language contamination in multilingual pre-training.
@@
In the training of models across various scales, we selected hyper-parameters based on experience and trial and error. Our preliminary results showed that deviating from optimal hyper-parameters can significantly harm model optimization and disrupt the scaling laws. To maintain consistency, we selected a constant learning rate, optimizer, learning rate scheduler, and batch size that matched the scale of the model for different experiments. This approach is in line with the conclusions of previous studies. Future research should explore the finding of optimal hyper-parameters from the perspective of language-specific scaling laws, which could lead to more effective pre-training configurations.
@@
Due to computational limitations, we were unable to cover a wide range of experiments, particularly in cases where the training data was extensive or the model size was very large. This limitation may reduce the generalizability of our findings to scenarios involving larger-scale models or datasets. In this study, we focused exclusively on the LLaMA2 architecture, which is recognized as a practical and effective transformer architecture for measuring scaling properties in pre-training. However, it is important to note that different architectures may have distinct scaling behaviors. This variability is a critical area for future investigation, as understanding these differences could provide deeper insights into optimizing and scaling various model architectures.

### v2

@@ 1 Introduction
Our main focus lies in the comparative analysis between the two strategies, including their scaling behaviors, the robustness of scaling, and their corresponding impact factors. We also study the technique of data replaying Ibrahim et al. (2024a) to mitigate catastrophic forgetting in CPT. Data replaying involves replaying a portion of the source language data during the training of the target language model. Finally, to explain and model the scalings under different settings, we fit a new extended scaling law for CPT, derived from Hoffmann et al. (2022). Our findings are outlined as follows:
@@ 2.1 Training Framework
To simplify the experiments, we use identical training strategies for all three pre-training sets, including learning rate schedules, batch sizes and trained token counts. All models are pre-trained with a context length of 2048 and undergo training on tokens equivalent to 20 times the model size (e.g., a 5B model is trained on 100B tokens). Although this is far from the extensive pre-training seen in recent practices  Touvron et al. (2023), as outlined in Hoffmann et al. (2022), the 20x trained token count is sufficient to demonstrate the loss-data scaling relationship. Our learning rate (LR) schedule features a cosine LR decay from a maximum LR of $2\times 10^{-4}$ and an LR warm-up, which increases the LR to the maximum in the first 5% of the training session. We use different batch sizes for different parameter sizes, as shown in Table 1.
@@ 2.2 Model and Data
Tokenizer
The tokenizer from LLaMA2 doesn’t properly represent common Chinese characters and tends to over-slice Chinese sentences. To prevent this, our tokenizer is trained on the bilingual pre-training corpus we used for the CPT experiments. The tokenizer is trained with SentencePiece Kudo (2018) and has a vocabulary size of 36,152, including common Chinese word pieces.
@@
Our English training data is primarily sampled from the RedPajama dataset Computer (2023), while the Chinese training data was sampled from Common Crawl Common Crawl (2007) and WuDao Xue et al. (2022). The raw text undergoes filtering and deduplication processes, which is similar to RedPajama Computer (2023). To study langauge robustness of the CPT strategy, we also conduct experiments on other languages, including French and Russian. We take their corresponding subsets from mC4 Raffel et al. (2019) as pre-training data. An total of $10^{6}$ tokens are held out from each respective training set as validation sets, remaining consistent across different models.
@@ 4.4 CPT Models Generalize to Downstream Tasks
Our results showed that in all three languages tested, the models enhanced through CPT consistently outperformed those trained from scratch, demonstrating improved performance across various languages and benchmarks.
@@ 5.1 What is the Compute-Optimal Allocation between Parameter Size and Data?
According to the framework established in Section 3, we can determine the optimal allocation between the model parameters $N_{opt}$ and training data $D_{opt}$ by minimizing the predicted loss $L$ with respect to data $D$ and parameter size $N$. More specifically, by optimizing Equation 2, we estimate the optimal training data and model parameters for pre-training from scratch to be:
@@
A visualization of the efficient frontier of model parameter $N$ with respect to compute over the IsoLoss contour is shown in Figure 5. We find that the optimal parameters for continual pre-training differ from those for pre-training from scratch, favoring less compute for the same model sizes. This aligns with the nature of cross-lingual transfer learning, where the model in continual pre-training is "pre-matured" due to prior knowledge acquired in the source language. This suggests that, in continual pre-training, using a larger language model is preferred over training on a larger dataset.
@@ 5.2 Does Replaying from Source Language Prevent Catastrophic Forgetting?
To prevent catastrophic forgetting of the original distributions during continual pre-training, we investigate methods that replay data from the source language during pre-training. We use the term replaying to refer to the practice of mixing data from the source language during continual pre-training on the target language. Previous works have shown that replaying data can help prevent catastrophic forgetting in continual learning tasks Ibrahim et al. (2024a); Scialom et al. (2022).
@@ 6 Related Work
Scaling laws help us understand how model performance changes with the amount of data and the size of the model. Kaplan et al. (2020) first introduced a detailed scaling law for large language models, demonstrating a clear relationship between model size, training data, and performance. Hoffmann et al. (2022) further explored this by emphasizing the trade-off between the data quantity and the model size, suggesting a compute-optimal allocation of data and parameters. Recent studies have examined scaling laws under specific conditions. Hernandez et al. (2022) and Muennighoff et al. (2023) focused on the diminishing returns from repeated tokens and excessive parameters. Tay et al. (2022) and Frantar et al. (2023) investigated how different model architectures impact scaling. Scaling laws are also relevant in the context of newer pre-training methods, such as parameter-efficient fine-tuning (PEFT) Kalajdzievski (2024) and Mixture-of-Experts (MoE) Krajewski et al. (2024).
@@
Transfer learning aims to enhance performance on new tasks by adapting pre-trained models with out-of-domain data. This process is more efficient when the source and target domains are closely related Pan and Yang (2009); Zhuang et al. (2020). Cross-lingual pre-training leverages language-independent knowledge embedded in pre-trained LLMs to improve performance in the target language Wu et al. (2019); Yosinski et al. (2014). Transfer learning is often studied within the context of limited-scale post-training, but it has been shown to be effective at a large pre-training scale with proper techniques Gupta et al. (2023). A significant challenge in transfer learning is catastrophic forgetting Winata et al. (2023), where the model’s ability in the original training domain degrades during transfer learning. Various strategies have been proposed to mitigate catastrophic forgetting, including modified learning rate schedules Ibrahim et al. (2024b); Gupta et al. (2023); Winata et al. (2023), data replay Ostapenko et al. (2022), and regularization Farajtabar et al. (2020). Our work combines data replay and modified learning rate schedules to combat catastrophic forgetting.
@@ 7 Conclusion
We used public datasets for pre-training, but completely preventing English contamination is challenging, especially since languages like French often include English words. Counting samples in each language split to estimate computational effort may be imprecise if other languages are present. Future research should analyze the impact of language contamination in multilingual pre-training more deeply.
@@
We selected hyper-parameters based on experience and trial and error when training models of various scales. Deviating from optimal hyper-parameters can significantly harm optimization and disrupt scaling laws. To maintain consistency, we used constant hyper-parameters matching the model scale, aligning with previous studies. Future research should find optimal hyper-parameters from the perspective of language-specific scaling laws for more effective pre-training.
@@
Due to computational limitations, we couldn’t conduct extensive experiments with large datasets or very large models, which may limit the generalizability of our findings to larger-scale scenarios. We focused on the LLaMA2 architecture, known for its practicality in measuring scaling properties. However, different architectures may exhibit distinct scaling behaviors. Future research should investigate these differences to better optimize and scale various model architectures.
Vocabulary Extension
We did not test the impact of vocabulary extension during continual pre-training (CPT). Instead, we used a byte-level BPE tokenizer trained on both English and Chinese text, keeping the model’s shape unchanged. While this allowed us to focus on scaling properties, it does not reflect practical scenarios where extending the vocabulary to include new tokens is necessary. This omission limits the applicability of our findings to cases where the tokenizer can not handle the new language well. Future work should explore vocabulary extension effects in CPT to provide a more comprehensive understanding.
Pre-training Length Variations
Our study assumed source models were trained to the Chinchilla-optimal token count. In practice, models are often "over-trained" beyond this point, such as LLaMA2 and LLaMA3 trained with trillions of tokens. We did not investigate how continual pre-training scales for these over-trained models. This limits the applicability of our scaling laws to real-world scenarios where models exceed the Chinchilla-optimal training length. Future research should examine CPT scaling for over-trained models to determine if our conclusions hold in such settings.
Acknowledgements
We would like to extend our special thanks to Yadong Liu and Chunhui Liu for their invaluable feedback and support.